A screw Jack who's pitch is used to raise a motor car of mass 900kg through a height of 20cm. The length of the Tommy bar of the Jack is 40cm. If the Jack is 60% efficient, calculate the
a.) work done by the effort (g= 10m\s)
b.) effort required
C.) mechanical advantage
d.) velocity ratio
1
Expert's answer
2018-11-05T15:42:09-0500

A=P·h=mgh=900·10·0.2=1800 J

M=Fl=(A·l)/s=(1800·0.4)/0.2=3600 N·m